Joseph leans a ladder that us 15 feet long against a building. The ladder is 12 feet away from the building. How far up does the ladder touch the side of the building?
Answer by mananth(16946)

The ladder ,ground and the wall form a right triangle
ground distance = leg1
height at which ladder touches
Hypotenuse = ladder length
Pythagoras theorem states that

Hypotenuse ^2= leg1^2+leg2^2

Hypotenuse = 15 ft
leg1 = 12 ft

leg2^2= hypotenuse ^2 - leg1^2

Leg2^2= 15^2-12 ^2
Leg2^2= 225 - 144
Leg2^2= 81
take the square root
leg2= 9 ft

height at which ladder touches =9 ft
m.ananth@hotmail.ca